OpenStudy (izhar):

3^a = 7^b = 21^c, express a in terms of b and c Can someone pls help me

OpenStudy (mww):

use logarithms. you will need to deal with them separately. You take a log of both sides (doesn't matter what log base you use, say base 10 or base e) Then use the rule log(a^n) = n log(a) - so the power gets taken outside of the log \[3^a = 7^b\] \[\log(3^a) = \log(7^b)\] \[a \log(3) = b \log(7)\] \[a = \frac{ b \log(7) }{ \log(3) }\] Try this for 3^a = 21^c

OpenStudy (izhar):

I get the idea but how can i get d solution from the question. The answer given for this question is a=(bc/ b-c)